How would you prepare 250.0 mL of 0.00200 M Na2S2O3? Describe what you would do in the lab. Include amounts and types of glassware and equipment that you would use.
Answer #1

Hi,


SAMPLE CALCULATION
To prepare the solution it is necessary to know the grams of Na2S2O3 that must be weighed, this is obtained by knowing the moles and subsequently with the molecular mass the grams.

The moles are known by multiplying the molarity (M) by the volume of the solution to be prepared in liters, we know the milliliters (250) but the liters are given as follows:

L = 250 ml * (1L / 1000ml) = 0.25 L

The moles then:

Moles = M * L
Moles = 0.002 mol / L * 0.25 L = 0.0005 mol

The molar mass of Na2S2O3 is set at: 158.11 g / mol

grams = moles * molar mass
grams = 0.0005 mol * 158.11 g / mol
grams = 0.08 g

EQUIPMENT, MATERIALS AND SUBSTANCES
- Watch glass
- Spatula
- Volumetric ball
- Graduated cylinder 250 ml
- Analytical balance

- Distilled water
- Na2S2O3

PREPARATION
- Measure 250 ml of distilled water with a graduated cylinder.
- Weigh 0.08 grams of Na2S2O3 on a watch glass.
- With the help of a spatula place the Na2S2O3 in a volumetric ball and finally add the distilled water before measured with the cylinder.
- Finally, stir to dissolve the solute.


If you liked the resolution, a thumbs up would help me a lot. Thank you very much and success. Regards